### v1

# A Survey on Efficient Federated Learning Methods for Foundation Model Training

Abstract
Federated Learning (FL) has become an established technique to facilitate privacy-preserving collaborative training. However, new approaches to FL often discuss their contributions involving small deep-learning models only. With the tremendous success of transformer models, the following question arises: What is necessary to operationalize foundation models in an FL application? Knowing that computation and communication often take up similar amounts of time in FL, we introduce a novel taxonomy focused on computational and communication efficiency methods in FL applications. This said, these methods aim to optimize the training time and reduce communication between clients and the server. We also look at the current state of widely used FL frameworks and discuss future research potentials based on existing approaches in FL research and beyond.
{textblock*}
5cm(1.52cm,26cm) Preprint. Under Review.
1 Introduction
Transformers, Foundation Models (FMs), and Large Langauge Models have conquered the deep learning world with unprecedented speed, enabling generative artificial intelligence for a broad audience. The most prominent example is OpenAI’s ChatGPT OpenAI (2023). Yet, these models require a carefully curated training dataset with several billion, or even trillion, samples. This limits the number of organizations capable of pre-training FMs to a handful of enterprises Zhou et al. (2023). Yet, fine-tuning methods for FMs have been introduced to leverage pre-trained FMs for their specific purposes (downstream tasks), which only require a fraction of the amount of data necessary for pre-training Kalyan et al. (2021). However, in practice, access to data is oftentimes challenging due to privacy regulations or environmental circumstances, such as poor connectivity between cloud and network edge devices Osia et al. (2020); Shokri and Shmatikov (2015).
To enable learning from massively distributed data and poorly connected clients with $<1\mathrm{GBit/s}$ bandwidth, McMahan et al. [2017] introduced Federated Learning (FL), a privacy-preserving machine learning technique that does not require the transfer of input data to a central server. However, this often comes at the cost of having non-identically and non-independently distributed (non-IID) data, variable client computational resources, and unstable communication, up to a point where the time spent on computation is as high as the time spent on communication Yousefpour et al. (2023). For FL applications, it is desirable to maximize computation and minimize communication. FMs, due to their large number of parameters Bommasani et al. (2021), already tend to take significantly longer to fine-tune than the models that are widely used in FL research to date He et al. (2020). As such, we see a natural focus on computation, which, in principle, can benefit us during training Ryabinin et al. (2023); Woisetschläger et al. (2023); Beutel et al. (2020). That said, it is more likely to use FL systems to fine-tune pre-trained FMs by, e.g., applying parameter-efficient fine-tuning techniques (PEFT) rather than pre-training. Pre-training generally requires full control over the training data and high data quality, which is hard to ensure in federated settings Longpre et al. (2023). PEFT often reduces the number of trainable parameters below 1% of all model parameters. As such, PEFT also helps reduce communication as we only need to transmit trainable parameters (see Section 3.2). This contrasts previously existing literature on efficiency methods for FL. There, computation and communication are perceived as separate topics, and experiments were carried out with small, full models.
Our survey is the first to study advances in computational and communication-efficient methods in FL applications with a special focus on FM training. Our work contains three distinct contributions:
  • A novel taxonomy on FMs in FL systems based on the core challenges in computation and communication. Our taxonomy focuses on computational and communication advancements that benefit FM training and fine-tuning.
  • In-depth discussion of current state computational and communication efficient methods for FL and FMs. FMs are large transformer models, often having several billions of parameters, making them expensive to communicate, train, or fine-tune. We discuss methods to facilitate efficient on-device training and fine-tuning. We also look into modern compression techniques for communicating large models.
  • A thorough overview of future research directions. We highlight future directions for research on computational and communication efficiency with regard to the harmonized evaluation of new methods. We further point out the increased complexity of hyperparameter optimization with PEFT and briefly touch upon the differences in privacy for full models vs. partial model updates for FMs.
Figure 1: Our Taxonomy. Foundation Models, in conjunction with Federated Learning, require efficient computational and communication methods.
2 Taxonomy
Our taxonomy introduces a novel perspective focused on the current developments in the field of efficient computational and communication methods for FL applications intended for training and fine-tuning large transformer models. While communication efficiency has been studied extensively in FL, computational advancements in conjunction with large models are currently emerging. Our taxonomy is visualized in Figure 1.
2.1 Basics of Federated Learning
To understand the relevancy of our taxonomy, it is key to briefly introduce the fundamentals of FL and the notations of this paper.
Coordinated by a server, the overall goal of FL is to collaboratively and iteratively train a DL model across a set of clients $N$ and minimize the global loss, where $f$ denotes an objective function with parameters $w$,
$\min_{w}\,f(w)\mathrel{\mathop{:}}=\frac{1}{|N|}\sum_{n=1}^{|N|}f_{n}(w)\mathrm{.}$ (1)
With this, we create a model that generalizes across all clients $n\in N$. Specifically, at the beginning, we commonly initialize the model weights across clients. Then, clients train the model on their local data and return the updated model parameters to the server. At the same time, each client updates its local parameters with fixed minibatch size $m$ over one or more epochs by applying gradient descent steps $\nabla l(w^{n}_{t};m)$ to the model,
$w_{t+1}^{n}=w_{t}-\eta\nabla l(w^{n}_{t};m),\,\forall n\in N\mathrm{.}$ (2)
Subsequently, only the model parameters $w_{t+1}^{n}$ are communicated back to the server.
2.2 Taxonomy Explanation
Our taxonomy is centered on two major challenges in FL: the computational and communication efficiency layers. With the emergence of FMs in FL applications, we see computational needs, as well as communication requirements, increase exponentially since models for real-world applications had a few 100K to 10M parameters in the past Wang et al. (2023); Huba et al. (2021). Now, models often reach billion-scale parameter sizes, i.e., a growth factor of more than 100×100$\times$100 × Zhang et al. (2023a).
Computational Efficiency. For smaller models, computation times on FL clients are seconds, while for large models, training can take several hours or even days. There are various techniques to lower computation times for FMs by reducing the number of trainable parameters.
Specifically, we discuss (I) full model training in FL systems as the longest-standing paradigm in FL (Section 3.1), (II) parameter-efficient fine-tuning as the most common technique for transformer fine-tuning (Section 3.2), (III) prompt tuning as a highly efficient method without training any model parameters (Section 3.3), and (IV) instruction tuning as a method to train a single model on heterogeneous downstream tasks (Section 3.4). For each technique, we outline its benefits and drawbacks, especially with regard to data heterogeneity.
Communication Efficiency. While computationally efficient methods for FM training in FL applications already reduce the number of parameters and data points to communicate, there is still a significant amount of data to be communicated between clients and servers. For instance, parameter-efficient fine-tuning methods only require 1–2% of parameters to be trainable. This still amounts to up to 14M parameters when working with Alpaca-7B Zhang et al. (2023a), larger than the majority of models currently being discussed in FL research He et al. (2020); Beutel et al. (2020).
We discuss communication efficiency methods along two major categories. (I) Model pruning is a method to communicate parts of a model between clients and the server, which resemble the most important parameters for a client (Section 4.1). (II) Full model compression is divided into three sub-categories: First, quantization is a method to decrease the numeric precision of model parameters. Second, sparsification is used as a way to zero out less important model parameters. Third, low-rank compression transforms high dimensional parameter matrices in client updates into significantly lower dimensional representations (Section 4.2).
Table 1: Computational Efficiency Methods for FL Systems and FM. Foundation Models (FMs) generally provide a strong performance across a variety of domains that are well explored in centralized learning Dosovitskiy et al. (2020); OpenAI (2023); Yang et al. (2023); Liu et al. (2023).
Training Underlying FL Max. Model Domain
Paper Regime Aggregation Strategy Params CV NLP Audio Graph
Centralized Learning FMT, PEFT, PT, IT $>100\mathrm{B}$
FedBERT Tian et al. (2022) FMT Weighted Average 117M
FedCLIP Lu et al. (2023) PEFT Weighted Average 85M
FedPEFT Sun et al. (2023) PEFT Weighted Average 85M
FedPETuning Zhang et al. (2023b) PEFT Weighted Average 125M
SLoRA Babakniya et al. (2023) PEFT Weighted Average 67M
FedPrompt Zhao et al. (2022) PT Weighted Average 223M
FedIT Zhang et al. (2023a) IT Weighted Average 7B
3 Federated Learning Methods for FMs
This section discusses recent methods to train FMs with FL methods. We distinguish between full model training (FMT), as it has been studied frequently in the FL domain, PEFT, prompt tuning (PT), and instruction tuning (IT). Table 1 summarizes existing computational efficiency methods for FM training.
3.1 Full Model Training
Generally, we refer to full model training when all parameters of a neural network are being trained. This is particularly useful when permanently fitting a model to a specific use case. In this approach, we locally train a model on all clients with the objective of minimizing $l$.
One of the first foundation model architectures is the BERT transformer. Tian et al. [2022] discuss federated pre-training of BERT-family models with up to 117M parameters for masked language modeling (MLM) and next-sentence prediction tasks. In MLM, the loss is defined over the sum of probabilities $P$ of predicted tokens $\hat{x}$ over the representation function of a masked sentence $g\big{(}\frac{x}{M(x)}\big{)}$,
$l(w^{t}_{n},m)=-\sum_{\hat{x}\in\mathcal{M(\mathrm{x})}}\mathrm{log}P\left(\hat{x}\left|g\left(\frac{x}{M(x)}\right)\right.\right),\,\forall n\in N\mathrm{.}$ (3)
MLM is the prevailing technique for self-supervised learning as it accelerates training on massive language bodies Raffel et al. (2019). Especially in federated environments, this enables one to learn from user data without the necessity of an expensive or manual labeling process.
Results show the effectiveness of federated pre-training and provide a net benefit over training on local datasets only Babakniya et al. (2023); He et al. (2020). However, the perplexity for large models (e.g., GPT-2) trained in a federated fashion is four orders of magnitude worse than centralized training, which is a stark indicator of low model quality Tian et al. (2022). However, for use cases involving sensitive data and strict privacy regulation, full model pre-training opens up the opportunity to create foundation models based on federated data.
Though full model training offers the highest degree of customization to a model, the technique requires substantial computational resources and communicating the entire model back and forth between clients and the server. In a federated setting, this can be very expensive and only feasible for a small number of participants. Therefore, more efficient methods are required.
3.2 Parameter Efficient Fine-Tuning
Generally, PEFT is used for large models that were trained on a large data basis already, provide good performance across a great variety of tasks, and require fine-tuning on a dataset for a specific downstream task to deliver excellent performance. This is especially effective since the data required for fine-tuning is orders of magnitude smaller than for pre-training. Also, since we use a pre-trained model and freeze the model parameters, the sensitivity towards non-IID data in fine-tuning is significantly lower than in pre-training. This renders PEFT techniques well-suited for FL applications and explains why this is a frequently discussed topic. Aside from the positive effects with regard to non-IID data, PEFT also reduces the computational costs for fitting a model to a specific downstream task. To achieve this, there are two methods as follows.
Sparse fine-tuning of pre-trained model parameters. As communication is a key concern in FL, it has become a priority to reduce the number of parameters to communicate between client and server. One of the most used approaches to achieve this is BitFit Zaken et al. (2021). The technique freezes almost all model parameters $w_{t}$ and only trains the bias term $b$ and the final classification layer $w^{\mathrm{final}}_{t}$ over the input features $a_{t}$, where the next-layer input features $a_{t+1}=a_{t}\cdot w_{t}+b$.
With this technique, it is only required to communicate $b$ and $w^{\mathrm{final}}_{t}$. Communication efforts are reduced by $\geq 99\%$, i.e., instead of communicating several 100M parameters, only a few 100K parameters are communicated.
Sun et al. [2023] propose FedPEFT, a framework for federated transformer fine-tuning that freezes the model weights to retain upstream knowledge within the model and adjust the systematic error for the downstream task. Their experimental results on vision transformers (ViT-B with 85M parameters) show on-par performance compared to full model fine-tuning on non-IID data, all the while reducing communication by 99.8%.
Adapter-based fine-tuning of additionally added parameters. When aiming to maintain a pre-trained model while introducing task-specific knowledge, adapter-based fine-tuning techniques provide strong performance and on-par efficiency as compared to sparse fine-tuning techniques. Here, we introduce two adapter layers in each transformer block of a foundation model. The adapter layer $a_{t+1}$ is calculated based on a downstream projection $w_{t}^{\mathrm{down}}$ of input feature $a_{t}$ into a lower-dimensional space $w_{t}^{\mathrm{down}}\in\mathcal{R}^{d\times r}$ followed by an upstream projection $w_{t}^{\mathrm{up}}\in\mathcal{R}^{r\times u}$, resulting in
$a_{t+1}=w_{t}^{\mathrm{up}}\cdot h(w_{t}^{\mathrm{down}}\cdot a_{t})\mathrm{.}$ (4)
With this, we can fine-tune an FM over much fewer dimensions than when fully fine-tuning a model. This saves both computational resources and communication on a client.
An improvement with regard to computational and communication efficiency over additional adapters is low-rank adapters (LoRA) Hu et al. (2021). The technique uses a lower dimensional representation $\mathbf{A}\in\mathcal{R}^{r\times u}$ and $\mathbf{B}\in\mathcal{R}^{d\times r}$ for weight updates $w_{t+1}=w_{t}+\Delta w=w_{t}+\mathbf{B}\mathbf{A})$. $r\ll\min(d,u)$ casts the weight update matrices into a much lower dimensionality than in the original transformer module without the necessity of adding additional adapters, i.e., LoRA builds an adapter for existing parameters. However, as $\mathbf{A}$ is initialized randomly to a Gaussian distribution and $\mathbf{B}$ as a zero matrix, this works well for centralized settings with IID data Hu et al. (2021). For FL settings, this initialization method bears the risk of slowing down the fine-tuning process over non-IID data.
FedCLIP Lu et al. (2023) introduces a PEFT method with an adjusted FedAvg-based adapter aggregation technique. Their approach yields significant performance improvements over vanilla FedAvg and FedProx.
Zhang et al. [2023b] provide a systematic benchmark study on adapter-based fine-tuning methods in privacy-preserving FL systems. Their results show that fine-tuning with additional adapters and LoRA both yield the same benchmark results regarding model accuracy. However, LoRA requires 3×3$\times$3 × less communication than additional adapters.
However, both adapter-based methods yield a worse accuracy than full fine-tuning. With SLoRA, Babakniya et al. [2023] propose a method to optimize LoRA for non-IID FL settings. Their approach entails parameterizing the weight update based on $r$,
$w_{t+1}=w_{t}+\frac{\beta}{r}\mathbf{B}\mathbf{A}\mathrm{.}$ (5)
As $\beta\in r$, scaling the fraction helps control the weight update impact of a single client. Subsequently, this can be used to control the negative effects caused by non-IID data. To practically achieve this, Babakniya et al. [2023] make use of a two-stage process: First, they used singular vector decomposition on $\mathbf{A}$ and $\mathbf{B}$ to obtain a common initialization point for LoRA across all clients in an FL system. Second, the training is facilitated with the commonly initialized low-rank representations. Their approach achieves on-par performance with full model fine-tuning. However, they require a warmup time of approximately 100 FL rounds for stage 1, which can be very expensive in FL settings as clients are often unavailable consecutively for such a long time McMahan et al. (2017).
3.3 Prompt Tuning
An equivalently efficient method for tuning pre-trained models to a downstream task is prompt tuning. Here, we use binary sentiments subsequent to masked-language-modeling to achieve high-quality results Lester et al. (2021). As such, the model remains entirely frozen, and we only train for a very small number of tokens that are being prepended to each embedded input query to improve the output quality. In contrast to fine-tuning, this method does not interfere with the model architecture or parameters. Lester et al. [2021] show that the effects of prompt tuning on model performance in centralized settings become better with larger models, i.e., for large FMs, prompt tuning bears significant potential.
Specifically, in FL settings, for each client $n$ the likelihood $P$ for a desired output $\hat{x}$ is calculated over prepending trainable embedded prompts $x_{p}$ to each embedded input $x$. In the interactive training process, the prompt is optimized in such a way that it optimally resembles the local objective of a client,
$\max(P_{n}(\hat{x}|[x_{p};x])),\,\forall n\in N\mathrm{.}$ (6)
Zhao et al. [2022] introduce FedPrompt, a method to efficiently communicate federally generated prompts only and aggregate them such that the global model performance of a pre-trained model improves for a downstream task. Their experimental evaluation shows a general sensitivity of prompt tuning towards data heterogeneity as the model performance degrades by 5 – 10% for the 100M parameter BERT model compared to a centrally trained baseline. However, with RoBERTa Base (124M parameters), the sensitivity diminishes, and the FL results are on par with centralized training. The larger T5 Base model (223M parameters) follows this trend, showing that prompt tuning becomes more effective with larger model sizes Lester et al. (2021).
3.4 Instruction Tuning
All fine-tuning methods require refinement of the pre-trained Natural Language Processing (NLP) FM to work well on downstream tasks. However, for some use cases, this is necessary due to their application in fields involving highly sensitive and strictly regulated data or requiring a very high model performance.
This is where instruction tuning comes into play as a technique that uses high-quality data. For instance, GPT-4 OpenAI (2023) uses Reinforcement Learning with Human Feedback (RLHF), which provides a high level of control over the training process and helps increase model quality. This, in general, requires very precise instructions to achieve the desired results. With FedIT, Zhang et al. [2023a] study instruction tuning on LLaMA-7B in an FL application over heterogeneous client tasks, e.g., learning brainstorming and text summarization on different clients in a single system at the same time. Their results show that the additional context on a downstream task generated with federated instruction tuning provides net benefits over central training only, even in heterogeneous settings. However, it is important to note that these results were produced over a single dataset only, Dollybricks-15k.
3.5 Discussion
As more and more open-source FMs become available for unrestricted use (e.g., Alpaca, Falcon), it is unlikely that full model training will be a common use case since training an FM from scratch in a centralized setting is very challenging. Also, for very sensitive data, where privacy is a key concern, like healthcare records, providers are establishing regulatory schemes for researchers to access and work with the data already de Kok et al. (2023).
Therefore, we see a priority in improving upon PEFT for downstream tasks, for instance, by introducing new and enhancing existing algorithms to remove the currently required warm-up times to improve the performance of LoRA in non-IID data environments. Computer Vision (CV) applications may benefit from exploring prompt tuning for vision transformers Jia et al. (2022). Also, we find significant challenges for instruction tuning as data quality is a general issue in FL Longpre et al. (2023) and access to trained staff, as it is required for RLHF, is hardly available in a real-world federated setting.
Furthermore, PEFT also positively affects communication efficiency as it significantly reduces the number of trainable (and thus communicable) parameters. As such, we now see the computational and communication efficiency grow closer together for FL and FMs.
Table 2: Communication-efficient FL methods. Their centralized learning pendants are often tied to specific domains: CV Habib et al. (2023), NLP He et al. (2021), Audio Perez et al. (2020), and Graph data Yang et al. (2021).
Enhancement Underlying FL Max. Model Domain
Algorithm Method(s) Aggregation Strategy Size (Params) CV NLP Audio Graph
Centralized Learning MP, Q, S, LRC $>100\mathrm{B}$
FedOBD Chen et al. (2023) Q Weighted Average 17M
PruneFL Jiang et al. (2022) MP Weighted Average 132M
FedPM Isik et al. (2022) MP Weighted Average 12M
FedTiny Huang et al. (2022) MP Weighted Average 132M
SoteriaFL Li et al. (2022) S FedSGD 0.05M
FjORD Horváth et al. (2021) MP Weighted Average 11M
H-FL Yang (2021) LRC Weighted Average 138M
FedPAQ Reisizadeh et al. (2020) Q FedSGD 0.09M
HeteroFL Diao et al. (2020) MP Weighted Average 11M
LotteryFL Li et al. (2020) MP Weighted Average 138M
4 Importance of Communication for FMs
With FMs, models have exponentially grown from a few million to several billion parameters to be able to resemble complex tasks. For FL, this specifically means that the communication load between clients and servers has grown significantly, even though with PEFT, there is not necessarily the need to communicate an entire model. However, when fine-tuning a multi-billion parameter FM with adapter-based methods, we still need to facilitate communication for more than a million parameters. For cross-device scenarios involving more than 1,000 clients per training round, the data traffic can quickly overstrain a server’s network capacity and potentially incur significant communication costs for data transfer from the edge to a cloud Xu et al. (2023); Isenko et al. (2023). Therefore, efficient communication and training design is essential for FL systems handling FMs. We distinguish two major efficiency methods: model pruning and full model compression. A detailed overview of studies on efficient communication in FL is provided in Table 2.
4.1 Model Pruning
The objective of model pruning (MP) is to retain and communicate only parts of a DL model that are relevant to a certain task. The reduction of parameters with this technique reduces the communication effort and the computational intensity Zhu and Gupta (2017). However, the success of pruning highly depends on the underlying data. Pruning client models without coordination may deny convergence with heterogeneous non-IID data in FL systems. Jiang et al. [2022] introduce PruneFL, a two-stage procedure to realize model pruning in FL systems. The first stage is carried out on a powerful client to find a common initialization for the model and generate the importance-based pruning mask. This mask is then iteratively refined over multiple FL training rounds under the consideration of all clients. The experimental results with PruneFL show two remarkable results: (I) The models have a shorter time to accuracy over the same task with PruneFL than with FedAvg, which is attributable to the higher degree of model specialization. (II) Since the model size is reduced, one would expect additional effects on faster computation, but there is limited hardware support for sparse matrix multiplications in training. Therefore, there are no computational benefits of PruneFL to this point. However, this may change with new hardware, such as sparse Tensor cores Zhu et al. (2019). With FedTiny, Huang et al. [2022] present an approach that works identically to PruneFL, except for them swapping the common initialization procedure with using batch normalization values of clients to choose a common initialization. Furthermore, FjORD Horváth et al. (2021) and HeteroFL Diao et al. (2020) provide similar approaches to pruning.
Isik et al. [2022] choose a similar approach for pruning models based on the lottery ticket hypothesis, first introduced to FL by Li et al. [2020]. Instead of commonly initializing a pruned model for an FL system, they initialize a random binary mask based on a shared seed on each client. This reduces computational efforts in the ramp-up phase. After an FL training round, each client communicates their binary mask to the server, which creates a global model based on the weighted average of those binary masks. With this, an approximate weight estimate replaces the parameters on the global model. From client to server, FedPM achieves significant communication efficiencies. However, the full model still has to be communicated from the server to the clients, lowering the net benefit.
Model pruning has also been discussed extensively for FM fine-tuning outside of FL Lagunas et al. (2021); Sanh et al. (2020). As existing pruning approaches have shown strong benefits to delivering on-par model performance compared to fine-tuning the full model, this is a promising direction to combine federated PEFT with highly efficient pruning techniques to further enhance communication. Along with pruning, sparse tensor computation hardware can also lower computational loads.
4.2 Full Model Compression
Model pruning is prone to omit segments of a DL model that may become relevant at a later stage. This originates from domain shifts Peng et al. (2020) and might require preserving the full model with all its parameters. For this, there are three frequently discussed techniques for full model compression in FL systems: Quantization, Sparsification, and Low-rank Compression.
Quantization (Q). The first work towards dynamic quantization is FedPAQ Reisizadeh et al. (2020), which combines FedAvg with strong quantization guarantees, where $Q$ represents the quantization term for a local model update,
$w_{t+1}=w_{t}+\frac{1}{|N|}\sum^{|N|}_{n=1}Q(w^{n}_{t+1}-w_{t})\mathrm{.}$ (7)
While DL models often operate on full precision (32-bit), this high degree of detail is not necessarily required Zhou et al. (2018). FedPAQ leverages this to reduce the communication intensity of FL applications. $Q$ calculates the optimal float precision of a model update to preserve all required information: $Q(w)=\mathinner{\!\left\lVert w\right\rVert}\cdot\mathrm{sign}(w)\cdot\xi(w,s)$, as proposed in QSGD by Alistarh et al. [2017]. $\xi$ formulates a stochastic process to dynamically tune $s$, the level of precision. FedPAQ has a significantly lower time to accuracy than QSGD, which is attributable to dynamizing $s$. However, it must be noted that dynamic quantization only yields benefits for communication. Depending on the infrastructure, the model updates have to be cast back to full precision, creating additional computational overhead on the client and server. Also, the method has been only tested with small models ($<100\mathrm{K}$ parameters).
FedOBD Chen et al. (2023) quantizes models with transformer block dropout, i.e., the random removal of entire model blocks. The dropout mechanism is carried out during training by each client and returns only the top-k most important model blocks to communicate. Additionally, FedOBD includes the ideas discussed in Alistarh et al. [2017] and Reisizadeh et al. [2020] but proposes an optimization problem out of the stochastic quantization where the trade-off originates from entropy and update size. The communication required for FL with a 17M parameter transformer model shows FedOBD to cut communication cost by 2×2$\times$2 × vs. FedPAQ and by 8×8$\times$8 × compared to vanilla FedAvg Chen et al. (2023).
Sparsification (S). While model pruning and sparsification technically have the same objective, pruned models do not necessarily resemble sparse models. A model is sparse once more than 50% of weights are set to 00 Frankle and Carbin (2019). However, pruning can also change the model architecture, i.e., not return the full model. Since it lends its idea from Frankle and Carbin (2019), FedPM Isik et al. (2022) (see section 4.1) can be considered as a model sparsification technique but does not necessarily lead to sparse networks and may return partial networks. SoteriaFL Li et al. (2022) guarantees sparse networks while maintaining differential privacy. Equation 7 is amended in such a way that $Q(w^{n}_{t+1}-w_{t})$ is replaced by $C(w^{n}_{t+1}+\mathcal{N}(0,\sigma^{2}\mathcal{I}))$ with $C$ resembling a sparse client update through shifted compression that has proven to improve convergence of DL models in FL settings compared to direct compression Mitra et al. (2021); Mishchenko et al. (2019). Overall, SoteriaFL mitigates the trade-off between model utility and compression, i.e., the differentially private models converge faster in stricter compression regimes than previously existing non-compressed differentially private approaches.
Low-rank Compression (LRC). H-FL Yang (2021) introduces a multi-level FL system with low-rank compression for full models by applying singular value decomposition to the feature matrix. This comes at the cost of model accuracy. To mitigate these effects, Yang [2021] includes a bias correction mechanism before each FL training round. However, it is to be noted that H-FL is a hierarchical learning framework that trains shallow models on the clients and employs mediating servers to connect the shallow client models with additional layers to form deep neural networks. The server is responsible for aggregating two models: the client models and the deep mediator network. Depending on the model architecture and nature of the task, maintaining two models could be cumbersome and incur additional communication overhead.
4.3 Discussion
To date, advancements in communication efficient methods for FL systems predominantly focus on training small full models. The communication paradigm shifts with the emergence of FMs in FL applications. With fine-tuning tasks, we only need to train a small fraction of model parameters, and thus, only trainable parameters have to be communicated. However, each trainable parameter usually contains a high degree of information for a downstream task. As such, the effectiveness of model pruning techniques is unclear as they would cut away fine-tuned parameters. At the same time, full-model compression techniques do not alter the model structure but rather reduce the parameter precision. Thus, these techniques can be used with FL applications with FMs in order to further reduce the size of the communicated updates. However, since fine-tuning is used for learning specific details in a downstream task, it is open to what degree we can compress PEFT model updates without removing essential information. Since PEFT is already sensitive to non-IID, efficient communication techniques can cause compounding effects as they remove information from trainable parameters.
Table 3: Current capabilities of state-of-the-art FL framework with respect to our taxonomy and their ability to run in resource-limited environments. The table is sorted by the frameworks’ ability to train FMs. Key: DP = Differential Privacy, HEC = Homomorphic Encryption, SMPC = Secure Multi-Party Computation.
Secure Training Communication FM Training / Edge
Framework Aggregation Efficiency Efficiency Fine-Tuning Ready
Clara DP, HEC
FedML DP, HEC PEFT
FederatedScope PEFT
Flower DP, SMPC PEFT
FATE HEC, SMPC PEFT
Substra
PySyft DP
OpenFL
TFF DP
IBM FL HEC
5 FMs in current FL frameworks
The backbones for making FL applications available to a broad audience are FL frameworks that implement recent advancements in FL research. Therefore, we investigate widely used frameworks for their FM readiness and capability with regard to computational and communication efficiency (Table 3).
Generally, all FL frameworks could handle FMs. Yet, only half implement efficient training methods. To further drive the adoption of FL in times of FMs, the frameworks need to improve both computationally and communication-efficient methods for training.
Clara is a proprietary FL framework designed specifically for medical applications and to work with NVIDIA GPUs NVIDIA (2021). The focus on powerful infrastructure makes training FMs in an FL setting possible. However, the Clara SDK does not include FL-specific optimizations to reduce computational or communication efforts specific to FMs. The final set of FL frameworks, characterized by their active open-source community, FATE Fan et al. (2023); Liu et al. (2021), FedML He et al. (2020), TensorFlow Federated (TFF) Google (2019), FederatedScope Kuang et al. (2023); Xie et al. (2023) and Flower Beutel et al. (2020), have adopted recent advancements in FL research. The frameworks especially allow for PEFT of FMs with LoRA.
Substra Galtier and Marini (2019), PySyft Ziller et al. (2021), OpenFL Foley et al. (2022), and IBM FL Ludwig et al. (2020), in their versions as of 2023, focus on training smaller FL tasks with several hundred thousand up to a few million parameters and, therefore, do not provide adapters for FM workloads with more than 100M parameters.
Yet, a consistent observation across all frameworks is their lack of efficient communication techniques. Workloads with FMs will significantly increase communication costs, and the growing use cases involving resource-constrained edge and IoT devices require high efficiency for computation and communication.
6 Related Work
While there are ample surveys that provide a broad perspective on FL Banabilah et al. (2022); Liu et al. (2022); Nguyen et al. (2021); Zhang et al. (2021), there are two closely related surveys to our work as they also focus on FMs.
In their survey, Zhuang et al. [2023] introduce a broad and general perspective on FMs and FL. They extensively discuss data modalities. This includes access to data across a large number of highly distributed clients and the quality of data that lives on these clients. Currently, FM training or fine-tuning requires datasets with high data quality, i.e., the instructions or texts used for MLM must be curated very carefully. Thus, their survey identifies a stark need for methods to train or fine-tune FMs on a scattered data basis with (highly) varying data quality. Further, Zhuang et al. [2023] discuss approaches to integrate FL applications into the lifecycle of FMs, i.e., how FMs can benefit from a continuously evolving system. While their survey briefly touches upon computational efficiency, our study provides an in-depth overview of state-of-the-art training and fine-tuning techniques to render FMs in FL applications a reality. Furthermore, our study includes a comprehensive overview of communication techniques that can enhance the adoption of FMs in communication resource-constrained environments.
Yu et al. [2023] provide an overview of FMs and FL with a special focus on privacy, an integral component of FL. Their survey includes a comprehensive overview of different fields of application for FMs, which they divide into once-off training and continual learning. The authors elaborate on technical challenges that may arise for specific use cases, such as robustness towards unreliable clients, varying data quality, the degree of non-IID data, and scalability. In contrast, our survey provides an application-agnostic, in-depth study of existing methods suitable for FM training. Our focus is to outline the technical challenges that currently hinder the operationalization of FM for use cases in federated applications.
7 Conclusions & Future Directions
In this paper, we survey the current landscape of computational and communication efficiency methods in FL systems and introduce a novel taxonomy based on the key techniques. While computational and communication efficiency have been separate topics in the past, they become closely intertwined as we start using FL systems to train FMs. Adjacent to our analysis of existing computational and communication efficiency methods in FL, we find the following three open questions that require attention from the FL community.
What are good and realistic evaluation strategies for generative downstream tasks in FL settings where we do not have control of data? Existing works focus on classification tasks. With the emergence of generative downstream tasks in FL systems, we require benchmarking tools that allow us to holistically validate the quality and efficiency of FL methods for FMs. It may be of particular interest to understand the notion of data heterogeneity in the context of generative tasks, i.e., how we define heterogeneous data splits.
How does hyperparameter optimization work for FMs in continuously evolving FL systems? While hyperparameter optimization in FL has been a key challenge, PEFT adds additional complexity. By the example of LoRA, we can systematically search for the optimal parameterization of $r$ while training in a centralized setting. In continuously evolving FL systems, parameterization needs to evolve over time. Furthermore, PEFT has been shown to be particularly sensitive to non-IID data, which further increases the complexity of hyperparameter optimization. SLoRA, for instance, shows that PEFT and FL can create good quality models, but to date, at significant costs in the form of long warm-up times.
How do PEFT and privacy in FL work together? Communication-efficient FL techniques, such as FedPM have been studied for their effect on privacy, but this is still an open topic for PEFT, PT, and IT. While it is proven that PEFT is more sensitive to data heterogeneity, the effects of perturbation through differential privacy are still subject to further studies. The same is true for PT and IT, as both techniques require precise prompts and instructions, respectively. As such, noise may have significantly negative effects here, as well.
Acknowledgments
This work is partially supported by the Bavarian Ministry of Economic Affairs, Regional Development and Energy (Grant: DIK0446/01).